# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v11

@@ C.1 Mine Deposit
Notes: These figures show how to trim the municipalities into the analytical samples. Figures C.1a and C.1b show the trimming steps for the coal and GSC (gold, silver, and copper) mines, respectively. Sources: Created by the author.
@@
Figure C.2: Location of Coal and GSC mines in 1931 and 1936
@@
Notes: The red and blue circles indicate the locations of coal and GSC (gold, silver, and copper) mines, respectively. Chikuhō coalfield in Fukuoka prefecture in Kyushū region is indicated with an arrow in Figure C.2a. Figure 3 illustrates the coal mines around Chikuhō coalfield in detail. Source: Created by the author.
@@ C.2 Geological Stratum
Note: The white circles indicate the location of the mines. The strata are illustrated through colors. Sources: Created by the author. Data on the location of mines are obtained from the Association for Harmonious Cooperation (1932, 1937). Data on the geological stratum are obtained from the Ministry of Land, Infrastructure, Transport and Tourism.
Figure C.3: Spatial distribution of mines and geological strata
@@
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator (Horn et al. 1975) are reported in parentheses. Notes: The number of observations for 1931 and 1936 are $1,140$ and $1,364$, respectively.
@@
Note: This geological columnar section shows the strata around the Iizuka coal mine, including coal layers. The first to third columns include the names of layers created in the Paleogene period. The fourth to seventh columns examine the length of each layer, columnar, and legend. The black layers shown in the columnar of the sixth column are the coal layers: more than ten coal layers are included in the layers created in the Paleogene period. Source: Fumoto (1961, p. Chishitsu chūjyō zu).
Figure C.4: Geological Columnar Section of the Iizuka Coal Mine
@@
Notes: This table shows the summary statistics for the indicator variable for the treatment group, geological stratum, and number of miners. The treatment (control) group is 0–5 (5-30) km from the centroid of a municipality that has a mine. The geological stratum is a time-constant variable: “Year” in parenthesis indicates the matched year. GSC mines indicate gold, silver, and copper mines. Sources: Data on the location of mines are from the Association for Harmonious Cooperation (1932, 1937).
@@ C.3 Dependent Variables
Marital fertility rate 1930$\ddagger$ $67.67$ $11.22$ $1140$ $64.23$ $10.84$ $167$ $68.26$ $11.18$ $973$
1935$\ddagger$ $67.92$ $11.07$ $1364$ $65.56$ $9.77$ $177$ $68.27$ $11.21$ $1187$
@@
Fetal death ${}^{\star}$ 1933$\ddagger$ $43.51$ $27.41$ $1140$ $51.57$ $26.25$ $167$ $42.13$ $27.38$ $973$
@@
Child mortality ${}^{\star}$ 1933$\ddagger$ $13.40$ $9.58$ $1140$ $15.54$ $7.11$ $167$ $13.03$ $9.90$ $973$
@@
Mining ${}^{\star}$ 1930$\ddagger$ $2.75$ $9.12$ $1140$ $17.13$ $17.64$ $167$ $0.28$ $1.63$ $973$
     ${}^{\star}$ 1930$\ddagger$ $3.16$ $10.09$ $1140$ $19.32$ $19.17$ $167$ $0.39$ $2.00$ $973$
     ${}^{\star}$ 1930$\ddagger$ $1.83$ $6.92$ $1140$ $11.85$ $14.30$ $167$ $0.11$ $0.99$ $973$
@@
Notes: $\ddagger$ and $\dagger$ in Panels B–D indicate that the mean difference between treatment and control groups is statistically significant at the 1% and 5% levels, respectively. The treatment (control) group is 0–5 (5-30) km from the centroid of a municipality that has a mine. Panel A reports the summary statistics for the demographic outcomes: population, crude marriage rates, crude birth rates, and household size. The population is log-transformed in the regressions to deal with a skewed distribution in nature. Household size is the total number of people divided by the number of households. The crude marriage rate is the number of marriages per $1,000$ people. The crude birth rate is the number of live births per $1,000$ people. The marital fertility rate is the number of live births per $1,000$ females. The crude death rate is the number of male (female) deaths per $1,000$ males (females). Panel B reports the summary statistics for infant mortality, fetal death, and child mortality rates. The infant mortality rate is the number of infant deaths per $1,000$ live births. The fetal death rate is the number of fetal deaths per $1,000$ births. The child mortality rate is the number of deaths of children aged 1–4 per $1,000$ children aged 1–5. The variables with $\star$ have a set of censored observations; the results from the Tobit models are materially similar to the main results (not reported). Panel C reports the summary statistics for labor force participation rates and employment shares. Labor force participation is the number of workers per $100$ people. The employment share is the number of workers in each sector per $100$ workers. Sources: Data on the number of people and households are from the Statistics Bureau of the Cabinet (1935b, 1939). Data on the demographic variables are from the Statistics Bureau of the Cabinet (1932, 1938). Data on the labor force participation rates and employment shares are from the Statistics Bureau of the Cabinet (1933). Data on births, fetal, infant, and child deaths are from Aiikukai (1935); Social Welfare Bureau (1941).
@@ C.3.1 Population, Marriage, Fertility, and Mortality
The official 1930 and 1935 Population Censuses published by the Statistics Bureau of the Cabinet document the number of people (by sex) and households in municipalities. The number of marriages and live births used to calculate crude marriage and birth rates were obtained from the Municipal Vital Statistics of 1930 and 1935 published by the Statistics Bureau of the Cabinet. The data on the number of deaths by gender are from the same documents.
@@ C.4.1 Railway Stations
Notes: The blue circles indicate the location of railway stations. The base map used in this figure is Figure 2. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures. Source: Created by the author.
@@ C.4.2 Seaports
Notes: The blue circles indicate the location of seaports. The ports in Okinawa prefecture and a few ports in the islands far from the main island are not shown in the figures. The base map used in this figure is Figure 2. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures. Source: Created by the author.
@@ C.4.3 Rivers
Notes: The blue lines indicate the location of rivers. The rivers in Okinawa prefecture and a few points in the islands far from the main island are not shown in the figures. The base map used in this figure is Figure 2. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures. Source: Created by the author.
@@ C.4.4 Elevation
Notes: This table shows the summary statistics for the control variables by type of mines. GSC mines indicate gold, silver, and copper mines. Distances to the nearest neighboring station, port, and river are in kilometers. Elevation is in meters. Sources: See Sections C.4.1 to C.4.4 for the data on the distance and elevation variables. Data on the location of mines are from the Association for Harmonious Cooperation (1932, 1937).
@@ D.1 Attenuation due to Measurement Error
[…]
@@ D.2 Assumptions for the Identification
[…]
@@
[…]
@@ D.3 Alternative Variance Estimator
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the cluster-robust covariance matrix estimator are reported in parentheses. Standard errors are clustered at the 107 (124) counties in the 1933 (1938) sample. Notes: Panel A: The results for the population (log-transformed), crude marriage rate, and marital fertility rate are reported. Panel B: ‘Male’ and ‘Female’ indicate the male death rate and female death rate, respectively. Panels C-1 and C-2: The results for the 1933 and 1938 Vital Statistics samples are listed. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. The IMR, FDR, and CDR indicate the infant, fetal, and child mortality rates, respectively. Panel D: ‘Sex Ratio’ and ‘HH Size’ indicate the sex ratio (male/female) and average household size, respectively. Table C.4 lists summary statistics on the dependent variables. All the estimates listed in Panels A, B, and D are from the IV estimator. All the regressions include the city and town fixed effects, railway accessibility, port accessibility, river accessibility, and elevation.
@@ D.4 Gender-Biased Structure Shift: Evidence from the 1930 Census
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses. Notes: Panels A and B show the results for male and female worker samples, respectively, from the 1930 Population Census (Panel C of Table C.4). Column 1 shows the results for the labor force participation rates (%), whereas Columns 2–6 show the results for employment share in each industrial sector (%).
@@ D.6 Testing the Validity of Threshold
Notes: This figure shows the results from the expanded regressions of equation 2 that includes the five indicator variables for each 5km bin from a mine to 25 km away from the origin. 0–5 km distance from mines indicates the municipalities within 5 km of a mine. 5–10 (10–15, 15–20, 20–25) km distance from mines indicates the municipalities within $(5,10]$ ($(10,15],(15,20],(20,25]$) km of a mine. (20–30] km bin is used as a reference group. Figures D.1aD.1bD.1cD.1dD.1e, and D.1f show the results for the local population, marital fertility rate, female mortality rate, infant mortality rate, sex ratio, and average household size, respectively. The dots and solid lines with caps show the estimates and their 95 percent confidence intervals, respectively. The confidence intervals are calculated using the standard errors based on the heteroskedasticity-robust covariance matrix estimator. Source: Created by the author.
@@
Notes: This figure shows the results from the expanded regressions of equation 2 that include the five indicator variables for each 5km bin from a mine to 25 km away from the origin. 0–5 km distance from mines indicates the municipalities within 5 km of a mine. 5–10 (10–15, 15–20, 20–25) km distance from mines indicates the municipalities within $(5,10]$ ($(10,15],(15,20],(20,25]$) km of a mine. (20–30] km bin is used as a reference group. Figure D.2 summarizes the results for the infant mortality rate in Fukuoka prefecture and the other prefectures. The dots and solid lines with caps show the estimates and their 95 percent confidence intervals, respectively. The confidence intervals are calculated using the standard errors based on the heteroskedasticity-robust covariance matrix estimator. Source: Created by the author.
in Fukuoka Prefecture
@@
Notes: Figure D.3a shows the spatial distribution of all the treatment bins used in Figure D.2. Figure D.3b highlights the location of the 10-15 km bin in Figure D.3a. Figure D.3c shows the infant mortality rates in 1938. Source: Created by the author.
@@ D.7 Sensitivity to the Regional Heterogeneity in Endowment of Male Miners
I assess the influence of regional heterogeneity in female labor patterns. Nishinarita (1985) argues that, due to its geographical characteristics, coal mines in Hokkaidō (a northernmost island) tended to employ men from rural villages in the Tōhoku (northeastern) region of the main island as their main miners. This forced the coal mines in Hokkaidō to use a sloping-face haulage method instead of using ato-yama. As a result, female miners in the Hokkaido mines were mainly engaged in the coal selection process rather than working in pits. It is unlikely that this regional heterogeneity could affect the location of mines, because the location of coal mines depends on the presence or absence of coal-bearing strata (Section 5.1). In addition, the number of mines in Hokkaidō is relatively small, meaning that such regional heterogeneity in labor supply is less likely to disturb the estimates. Despite this, to test whether such features lead to omitted variable bias, I added an indicator variable for the municipalities of the Hokkaidō prefecture in the regressions. Table D.3 shows the results from the specification, including an indicator variable for the municipalities in the Hokkaidō prefecture. The results are materially similar to those of the main results presented in the main text.
@@
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses. Notes: Panel A: The results for the population (log-transformed), crude marriage rate, and marital fertility rate are reported. Panel B: ‘Male’ and ‘Female’ indicate the male death rate and female death rate, respectively. Panels C-1 and C-2: The results for the 1933 and 1938 Vital Statistics samples are listed. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. The IMR, FDR, and CDR indicate the infant, fetal, and child mortality rates, respectively. Panel D: ‘Sex Ratio’ and ‘HH Size’ indicate the sex ratio (male/female) and average household size, respectively. Table C.4 lists summary statistics on the dependent variables. All the estimates listed in Panels A, B, and D are from the IV estimator. All the regressions include the city and town fixed effects, railway accessibility, port accessibility, river accessibility, elevation, and a Hokkaidō dummy.

### v12

@@ C.1 Mine Deposit
Notes: These figures show how to trim the municipalities into the analytical samples. Figures C.1a and C.1b show the trimming steps for the coal and GSC (gold, silver, and copper) mines, respectively.
Sources: Created by the author.
@@
Notes: The red and blue circles indicate the locations of coal and GSC (gold, silver, and copper) mines, respectively. Chikuhō coalfield in Fukuoka prefecture in Kyushū region is indicated with an arrow in Figure C.2a. Figure 3 illustrates the coal mines around Chikuhō coalfield in detail.
Source: Created by the author.
@@ C.2 Geological Stratum
Note: The white circles indicate the location of the mines. The strata are illustrated through colors.
Sources: Created by the author. Data on the location of mines are obtained from the Association for Harmonious Cooperation (1932, 1937). Data on the geological stratum are obtained from the Ministry of Land, Infrastructure, Transport and Tourism.
@@
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator (Horn et al. 1975) are reported in parentheses.
Notes: The number of observations for 1931 and 1936 are $1,140$ and $1,364$, respectively.
@@
Note: This geological columnar section shows the strata around the Iizuka coal mine, including coal layers. The first to third columns include the names of layers created in the Paleogene period. The fourth to seventh columns examine the length of each layer, columnar, and legend. The black layers shown in the columnar of the sixth column are the coal layers: more than ten coal layers are included in the layers created in the Paleogene period.
Source: Fumoto (1961, p. Chishitsu chūjyō zu).
@@
Notes: This table shows the summary statistics for the indicator variable for the treatment group, geological stratum, and number of miners. The treatment (control) group is 0–5 (5-30) km from the centroid of a municipality that has a mine. The geological stratum is a time-constant variable: “Year” in parenthesis indicates the matched year. GSC mines indicate gold, silver, and copper mines.
Sources: Data on the location of mines are from the Association for Harmonious Cooperation (1932, 1937).
@@ C.3 Dependent Variables
Marital fertility rate 1930$\ddagger$ $174.14$ $29.69$ $1140$ $161.91$ $30.58$ $167$ $176.24$ $29.04$ $973$
1935$\ddagger$ $177.50$ $29.75$ $1364$ $169.40$ $27.87$ $177$ $178.70$ $29.84$ $1187$
@@
Fetal death rate 1933$\ddagger$ $43.51$ $27.41$ $1140$ $51.57$ $26.25$ $167$ $42.13$ $27.38$ $973$
@@
Child mortality rate 1933$\ddagger$ $13.40$ $9.58$ $1140$ $15.54$ $7.11$ $167$ $13.03$ $9.90$ $973$
@@
Mining sector 1930$\ddagger$ $2.75$ $9.12$ $1140$ $17.13$ $17.64$ $167$ $0.28$ $1.63$ $973$
       Male 1930$\ddagger$ $3.16$ $10.09$ $1140$ $19.32$ $19.17$ $167$ $0.39$ $2.00$ $973$
       Female 1930$\ddagger$ $1.83$ $6.92$ $1140$ $11.85$ $14.30$ $167$ $0.11$ $0.99$ $973$
@@
Notes: $\ddagger$ and $\dagger$ in Panels B–D indicate that the mean difference between treatment and control groups is statistically significant at the 1% and 5% levels, respectively. The treatment (control) group is 0–5 (5-30) km from the centroid of a municipality that has a mine.
Panel A reports the summary statistics for the demographic outcomes: population, crude marriage rates, crude birth rates, and household size. The population is log-transformed in the regressions to deal with a skewed distribution in nature. Household size is the total number of people divided by the number of households. The crude marriage rate is the number of marriages per $1,000$ people. The crude birth rate is the number of live births per $1,000$ people. The marital fertility rate is the number of live births per $1,000$ married females. The crude death rate is the number of male (female) deaths per $1,000$ males (females).
Panel B reports the summary statistics for infant mortality, fetal death, and child mortality rates. The infant mortality rate is the number of infant deaths per $1,000$ live births. The fetal death rate is the number of fetal deaths per $1,000$ births. The child mortality rate is the number of deaths of children aged 1–4 per $1,000$ children aged 1–5. The variables with $\star$ have a set of censored observations; the results from the Tobit models are materially similar to the main results (not reported).
Panel C reports the summary statistics for labor force participation rates and employment shares. Labor force participation is the number of workers per $100$ people. The employment share is the number of workers in each sector per $100$ workers.
Sources: Data on the number of people and households are from the Statistics Bureau of the Cabinet (1935b, 1939). Data on the demographic variables are from the Statistics Bureau of the Cabinet (1932, 1938). Data on the labor force participation rates and employment shares are from the Statistics Bureau of the Cabinet (1933). Data on births, fetal, infant, and child deaths are from Aiikukai (1935); Social Welfare Bureau (1941).
@@ C.3.1 Population, Marriage, Fertility, and Mortality
The official 1930 and 1935 Population Censuses published by the Statistics Bureau of the Cabinet document the number of people (by sex), households, and married females in municipalities. The number of marriages and live births used to calculate crude marriage and birth rates were obtained from the Municipal Vital Statistics of 1930 and 1935 published by the Statistics Bureau of the Cabinet. The data on the number of deaths by gender are from the same documents.
@@ C.4.1 Railway Stations
Notes: The blue circles indicate the location of railway stations. The base map used in this figure is Figure 2. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author.
@@ C.4.2 Seaports
Notes: The blue circles indicate the location of seaports. The ports in Okinawa prefecture and a few ports in the islands far from the main island are not shown in the figures. The base map used in this figure is Figure 2. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author.
@@ C.4.3 Rivers
Notes: The blue lines indicate the location of rivers. The rivers in Okinawa prefecture and a few points in the islands far from the main island are not shown in the figures. The base map used in this figure is Figure 2. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author.
@@ C.4.4 Elevation
Notes: This table shows the summary statistics for the control variables by type of mines. GSC mines indicate gold, silver, and copper mines. Distances to the nearest neighboring station, port, and river are in kilometers. Elevation is in meters.
Sources: See Sections C.4.1 to C.4.4 for the data on the distance and elevation variables. Data on the location of mines are from the Association for Harmonious Cooperation (1932, 1937).
@@ D.1 Attenuation due to Measurement Error
Consider a simplified projection of $y_{i}$ on $\textit{MineDeposit}_{i}$ as $y_{i}=\theta\textit{MineDeposit}_{i}+u_{i}$. When $\textit{MineDeposit}_{i}$ has a classical measurement error, as $\textit{MineDeposit}^{*}_{i}=\textit{MineDeposit}_{i}+u_{i}$, the observable model becomes $y_{i}=\theta^{*}\textit{MineDeposit}^{*}_{i}+\nu_{i}$. Consequently, I use $\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})=E[\textit{MineDeposit}^{*}_{i}\nu]=-\theta\text{var}(u_{i})\neq 0$ for $\theta^{*}=\theta+\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})/\text{var}(\textit{MineDeposit}^{*}_{i})$.
@@ D.2 Assumptions for the Identification
$\footnotesize{\begin{split}y_{i}=\alpha+\beta\text{{MineDeposit}}_{i}+e_{i},\\ \text{{MineDeposit}}_{i}=\kappa+\tau\text{{Stratum}}_{i}+\epsilon_{i}.\end{split}}$ (20)
@@
Second, let $\textit{MineDeposit}_{i}(\textit{Stratum}_{i})$ be an indicator for whether $i$ is treated given the random assignments of the geological strata, under the SUTVA. The nonzero average causal effect indicates that $E[\textit{MineDeposit}_{i}(1)-\textit{MineDeposit}_{i}(0)]\neq 0$. This assumption clearly holds because the likelihood of having coal mines increases with the IV strata (Online Appendix C.2).
Third, the monotonicity assumption suggests that $\forall i$, $\textit{MineDeposit}_{i}(1)\geq\textit{MineDeposit}_{i}(0)$. In sum, a combination of the nonzero average causal effect and monotonicity assumptions argue that strong monotonicity should be satisfied at least in one unit as $\textit{MineDeposit}_{i}(1)>\textit{MineDeposit}_{i}(0)$ (Angrist et al. 1996). Although this assumption is an idea for the potential outcome and is untestable (Imbens and Angrist 1994), there is fundamentally no reason for mining companies to avoid the stratum that contains coal seams. Thus, the instrument should affect the selection decision monotonically.
@@ D.3 Alternative Variance Estimator
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the cluster-robust covariance matrix estimator are reported in parentheses. Standard errors are clustered at the 107 (124) counties in the 1933 (1938) sample.
Notes: Panel A: The results for the population (log-transformed), crude marriage rate, and marital fertility rate are reported. Panel B: ‘Male’ and ‘Female’ indicate the male death rate and female death rate, respectively. Panels C-1 and C-2: The results for the 1933 and 1938 Vital Statistics samples are listed. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. The IMR, FDR, and CDR indicate the infant, fetal, and child mortality rates, respectively. Panel D: ‘Sex Ratio’ and ‘HH Size’ indicate the sex ratio (male/female) and average household size, respectively. Table C.4 lists summary statistics on the dependent variables. All the estimates listed in Panels A, B, and D are from the IV estimator. All the regressions include the city and town fixed effects, railway accessibility, port accessibility, river accessibility, and elevation.
@@ D.4 Gender-Biased Structure Shift: Evidence from the 1930 Census
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Panels A and B show the results for male and female worker samples, respectively, from the 1930 Population Census (Panel C of Table C.4). Column 1 shows the results for the labor force participation rates (%), whereas Columns 2–6 show the results for employment share in each industrial sector (%).
@@ D.6 Testing the Validity of Threshold
Notes: This figure shows the results from the expanded regressions of equation 2 that includes the five indicator variables for each 5km bin from a mine to 25 km away from the origin. 0–5 km distance from mines indicates the municipalities within 5 km of a mine. 5–10 (10–15, 15–20, 20–25) km distance from mines indicates the municipalities within $(5,10]$ ($(10,15],(15,20],(20,25]$) km of a mine. (20–30] km bin is used as a reference group.
Figures D.1aD.1bD.1cD.1dD.1e, and D.1f show the results for the local population, marital fertility rate, female mortality rate, infant mortality rate, sex ratio, and average household size, respectively. The dots and solid lines with caps show the estimates and their 95 percent confidence intervals, respectively. The confidence intervals are calculated using the standard errors based on the heteroskedasticity-robust covariance matrix estimator.
Source: Created by the author.
@@
Notes: This figure shows the results from the expanded regressions of equation 2 that include the five indicator variables for each 5km bin from a mine to 25 km away from the origin. 0–5 km distance from mines indicates the municipalities within 5 km of a mine. 5–10 (10–15, 15–20, 20–25) km distance from mines indicates the municipalities within $(5,10]$ ($(10,15],(15,20],(20,25]$) km of a mine. (20–30] km bin is used as a reference group.
Figure D.2 summarizes the results for the infant mortality rate in Fukuoka prefecture and the other prefectures. The dots and solid lines with caps show the estimates and their 95 percent confidence intervals, respectively. The confidence intervals are calculated using the standard errors based on the heteroskedasticity-robust covariance matrix estimator.
Source: Created by the author.
@@
Notes: Figure D.3a shows the spatial distribution of all the treatment bins used in Figure D.2. Figure D.3b highlights the location of the 10-15 km bin in Figure D.3a. Figure D.3c shows the infant mortality rates in 1938.
Source: Created by the author.
@@ D.7 Sensitivity to the Regional Heterogeneity in Endowment of Male Miners
I assess the influence of regional heterogeneity in female labor patterns. Nishinarita (1985) argues that, due to its geographical characteristics, coal mines in Hokkaidō (a northernmost island) tended to employ men from rural villages in the Tōhoku (northeastern) region of the main island as their main miners (see also Ogino 1993, p. 32). This forced the coal mines in Hokkaidō to use a sloping-face haulage method instead of using ato-yama. As a result, female miners in the Hokkaido mines were mainly engaged in the coal selection process rather than working in pits. It is unlikely that this regional heterogeneity could affect the location of mines, because the location of coal mines depends on the presence or absence of coal-bearing strata (Section 5.1). In addition, the number of mines in Hokkaidō is relatively small, meaning that such regional heterogeneity in labor supply is less likely to disturb the estimates. Despite this, to test whether such features lead to omitted variable bias, I added an indicator variable for the municipalities of the Hokkaidō prefecture in the regressions. Table D.3 shows the results from the specification, including an indicator variable for the municipalities in the Hokkaidō prefecture. The results are materially similar to those of the main results presented in the main text.
@@
***, **, and * represent statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors based on the heteroskedasticity-robust covariance matrix estimator are reported in parentheses.
Notes: Panel A: The results for the population (log-transformed), crude marriage rate, and marital fertility rate are reported. Panel B: ‘Male’ and ‘Female’ indicate the male death rate and female death rate, respectively. Panels C-1 and C-2: The results for the 1933 and 1938 Vital Statistics samples are listed. Columns 1–2 show the results for the entire sample. Columns 3–4 show the results for the municipalities with coal mines (i.e., the municipalities included in the treatment group). Columns 5–6 show the results for the full sample and the municipalities with coal mines, respectively. The IMR, FDR, and CDR indicate the infant, fetal, and child mortality rates, respectively. Panel D: ‘Sex Ratio’ and ‘HH Size’ indicate the sex ratio (male/female) and average household size, respectively. Table C.4 lists summary statistics on the dependent variables. All the estimates listed in Panels A, B, and D are from the IV estimator. All the regressions include the city and town fixed effects, railway accessibility, port accessibility, river accessibility, elevation, and a Hokkaidō dummy.